A learner adds two fractions as follows 2/3 + 5/6 =7/9 define the problem which you have to work on with this learner
1
Expert's answer
2020-07-07T20:38:51-0400

The concept of LCM has to be understood first. When adding two fractions their respective denominators are used to get the LCM. In this scenario the denominators are 3 and 6. The LCM of 3 and 6 is 6. The LCM can be found using the table method or the factors method.

3 = 3 X 1

6 = 2 X 3

LCM = 1 X 2 X 3 = 6

The LCM is then put as the denominator of the answer, fractions are brought to the common denominator:

"2\/3 + 5\/6 = (4+5)\/6 = 9\/6"

9/6 is further simplified to 3/2 = 1 +1/2